Kate has 2 more apples than she needs in order to have one-third of the apples that Renee has. Kate is 8 apples short of having half of the apples that Renee has. How many apples does Kate have?
Found 2 solutions by stanbon, Aldorozos:
Answer by stanbon(75887)

Kate has 2 more apples than she needs in order to have one-third of the apples that Renee has. Kate is 8 apples short of having half of the apples that Renee has. How many apples does Kate have?
-----
Equations:
k-2 = (1/3)r
k+8 = (1/2)r
------------
r = 3k-6
r = 2k+16
-----
3k-6 = 2k+16
k = 22 apples (# Kate has)
----------------
Cheers,
Stan H.
-----------

Answer by Aldorozos(172)

k-2 = 1/3 R
k+8 = 1/2/R
Let's subtract the first equation from the second
k-k + 8-(-2) = 1/2 R - 1/3 r
60 = 1/6 R therefore R = 60 = number of apples Renee has
k+8 = 1/2 60 Therefore K =22
You may check the answer by subtracting 2 from 22 to see if it is 1/3 of 60